# Reducing leakage of single-qubit gates for superconducting quantum processors using analytical control pulse envelopes

### v1

@@ IV Outlook
The developed pulse shaping methods are not only limited to superconducting qubits or to the suppression of the $ef$-transition of a single transmon qubit, but they may also be applied to other platforms or to suppress multiple leakage transitions. Thus, the proposed methods may improve simultaneous single-qubit gates on quantum processing units of any type that suffer from frequency crowding issues. This may be achieved by minimizing the spectral energy of the control pulses across crosstalk-induced transitions in addition to the $ef$-transition, while using a cutoff frequency to regularize the pulse shape. The proposed methods may also find applications in the optimization of two-qubit gate control pulses based on flux pulses [5, 4] or microwave pulses [6]. Furthermore, we expect that combining the proposed methods with experimental closed-loop optimization may result in further improvements to the fidelity of fast quantum logic gates in systems with high coherence. The proposed methods provide good initial parameter guesses and a natural basis for the optimization, potentially reducing the required number of iterations and wall clock time significantly.
@@ Acknowledgements.
We thank the whole staff at IQM Quantum Computers for their support. Especially, we acknowledge Lucas Ortega, Roope Kokkoniemi and Matthew Sarsby for supporting the maintenance and improvement of the experimental setup; Tuure Orell, Jani Tuorila, and Hao Hsu for simulation-related discussions; and Attila Geresdi on discussions related to single-qubit gates. E.H. thanks the Finnish Foundation for Technology Promotion (grant No. 9230) and Jenny and Antti Wihuri Foundation (grant No. 00230115) for funding. We acknowledge the provision of facilities and technical support by Aalto University at OtaNano - Micronova Nanofabrication Center and LTL infrastructure. All authors declare that IQM has filed a patent application regarding the new methods for shaping the frequency spectrum of control pulses having the following inventors: Eric Hyyppä, Antti Vepsäläinen and Johannes Heinsoo. E.H. developed the concept and theory for FAST DRAG and HD DRAG. E.H., A.V., and J.H. planned the simulations and the experiments. E.H. conducted the simulations with support from M.P. and S.I.. E.H. conducted the experiments with support from J.L., F.M., S.O., and C.F.C. on measurement code, F.M. on experiment setup, and B.T. on experiment software. A.L. and C.O.-K designed the sample. W.L. fabricated the device. The manuscript was written by E.H. with support from A.V., and J.H.. All authors commented on the manuscript. A.V., and J.H. supervised the work.
@@ Appendix A Derivation of FAST pulse shapes
where $N$ is the number of basis functions, and $\{g_{n}(t)\}$ is the set of basis functions with associated coefficients $\{c_{n}\}$. In this work, we have chosen the sum of basis functions to represent a Fourier cosine series parametrized such that the control envelope $\Omega_{\mathrm{I}}:\mathbb{R}\rightarrow\mathbb{R}$ and its derivative have no discontinuities, thus reducing slow sinc-type decay of the pulse spectrum. Thus, we write the control envelope $\Omega_{\mathrm{I}}(t)$ as
$\Omega_{\mathrm{I}}(t)=A\times\sum_{n=1}^{N}c_{n}[1-\cos(2\pi nt/t_{\mathrm{p}})].$ (9)
@@
$=\int_{0}^{t_{\mathrm{g}}}[1-\cos(2\pi nt/t_{\mathrm{p}})]\mathrm{e}^{-\mathrm{i}2\pi tf}\mathrm{d}t$
@@
where $\tilde{\bm{c}}=(\bm{c}^{T},\mu)^{T}$ is the coefficient vector extended by the Lagrangian multiplier, $\tilde{\bm{b}}=(\bm{0}^{T},\theta/t_{\mathrm{p}})^{T}\in\mathbb{R}^{(N+1)\times 1}$, and the matrix $\tilde{\bm{A}}$ is defined as
@@
Subsequently, we briefly discuss some useful properties of the FAST method. First, we emphasize that the FAST method enables the construction of pulses with a given duration. Thus, the method differs from filter-based solutions, such as the use of band-pass or low-pass filters [16], which may filter out specific frequency intervals from the spectrum of a control pulse but which also result in pulse distortions, potentially significantly changing the pulse duration. As already discussed in Sec. II.2, the FAST method enables one to control the width and strength of the spectral suppression. Thus, a combination of FAST and DRAG may be helpful to reach a lower leakage error for fast logic gates compared to conventional DRAG pulses since the spectral suppression provided by conventional DRAG pulses may not be sufficient, e.g., in the presence of multiple leakage transitions or frequency shifts occurring during the gate. However, one should bear in mind that the suppression of several frequency intervals using the FAST method may lead to increased spectral power at non-suppressed frequencies.
@@ Appendix B Higher-derivative DRAG
$\Omega_{\mathrm{Q}}(t)$ $=-\frac{A\beta}{\alpha}\dot{\Omega}_{\mathrm{I}},$ (22)
@@
The formulation of HD DRAG in Eqs. (21) and (22) decouples the in-phase parameters $\{\beta_{2n}\}_{n=1}^{K}$ and associated suppressed frequencies $\{f_{j}\}_{j=1}^{K}$ from the DRAG coefficient $\beta$, which simplifies the calibration experiments especially for the case of $K=1$ considered in Sec. II.3 and also enables one to tune $\beta$ to avoid phase errors if desired (see Sec. II.1). By solving both $\{\beta_{2n}\}_{n=1}^{K}$ and $\beta$ to suppress a given undesired baseband frequency $f_{\mathrm{s}}$, a stronger and wider spectral suppression can be obtained than with conventional DRAG pulses. Furthermore, the in-phase parameters $\{\beta_{2n}\}_{n=1}^{K}$ do not affect the effective drive frequency (unlike $\beta$), which can beneficial for calibration experiments in some scenarios.
@@ Appendix C Single-qubit gate simulations
In this section, we describe the methods used to simulate the gate error and leakage as illustrated in Fig. 3, and in Supplementary Figs. 9--12. For the simulations, we model the bare transmon using an anharmonic Duffing oscillator Hamiltonian as
@@
For a given control pulse shape and gate duration, we set the drive frequency equal to the qubit frequency and calibrate the pulse parameters $\{A,\beta,\varphi_{z}\}$ for a $R_{X}(\pi/2)$-gate based on DRAG-P or DRAG-L using a simulated calibration sequence similar to the one used in the experiments and described in Appendix. F. As a minor difference, the leakage per gate needed for calibrating $\beta$ in the case of DRAG-L is estimated based on a single simulated $R_{X}(\pi/2)$-gate instead of using a leakage RB sequence.
@@ Appendix D Experimental setup
To implement single-qubit rotations, the qubits are capacitively coupled to individual drive lines that are connected to room-temperature electronics via 50 dB of total nominal attenuation separated across different temperature stages of the cryostat. At the qubit frequency, the total room-temperature attenuation in the cryostat was characterized to be approximately 67 dB. At room temperature, the single-qubit control pulses are generated at an intermediate frequency using a channel pair of a HDAWG (Zurich Instruments) with a sampling rate of 2.4 GSa/s and a peak-to-peak maximum amplitude range of 3 V or 4 V enabled by an internal amplifier, which limits the analog bandwidth to 300 MHz. To attain a sample-level granularity for the gate duration and facilitate envelope predistortion discussed in Ap. G, we evaluate the $I$ and $Q$ envelope waveforms of each gate sequence on a computer accounting for the Virtual Z-rotations, after which we send these long waveforms to the HDAWG. The intermediate-frequency control pulses are upconverted to the qubit frequency using an IQ mixer and an in-house-built local oscillator with an integrated root-mean-squared phase jitter below 4 mrad from 1 kHz to 10 MHz.
@@
For qubit state readout, the qubits are capacitively coupled to individual readout structures consisting of a readout resonator and a Purcell filter with both of the readout structures further coupled to a common probe transmission line. We generate the readout pulses at an intermediate frequency using an UHFQA (Zurich Instruments) and upconvert the pulses with an IQ mixer and a single SGS100A microwave generator (Rohde & Schwarz) that is used both for the up-conversion and down-conversion of the readout signal. On its way to the sample, the readout signal is passed through 60 dB of nominal attenuation. The readout signal transmitted through the probe line is amplified using an high-electron-mobility transistor (HEMT) at the 4 K stage with further amplification applied at room temperature. The amplified readout signal is downconverted back to an intermediate frequency and subsequently digitized using the UHFQA. Note that our setup does not include a quantum-limited parametric amplifier despite of which we reach a decent two-state readout assignment fidelity of up to $\sim 95$%.
@@ Appendix E Three-state readout
In this section, we briefly describe our approach for performing three-state readout on the qubit used for the single-qubit gate studies. At the flux-insensitive operation point, we measure a dispersive shift of $\chi/(2\pi)=4.0$ MHz. For the readout pulse, we use a readout frequency of 6.146 GHz and a relatively long pulse duration of 1.5 $\mu$s due to the absence of a quantum-limited amplifier in the readout chain. For the integration of the acquired readout signal, we use an integration time of 1.5 $\mu$s and set all integration weights to unity. To calibrate the classification boundaries for three-state readout, we repeatedly prepare the qubit in each of the states $|0\rangle$, $|1\rangle$, and $|2\rangle$, and record $326782^{15}=$ 32678 single-shot IQ voltages for each state as shown in Fig. 7. For each prepared state, the mean of the acquired IQ voltages is computed and for new readout events, the qubit is assigned to the state with the smallest distance between the recorded single-shot IQ voltage and the state mean. The resulting readout assignment matrix is given by
@@
Appendix F Single-qubit gate calibration

### v2

@@ IV Outlook
The developed pulse shaping methods are not only limited to superconducting qubits or to the suppression of the $ef$-transition of a single transmon qubit, but they may also be applied to other platforms or to suppress multiple leakage transitions or other undesired transitions. Thus, the proposed methods may improve simultaneous single-qubit gates on quantum processing units of any type that suffer from frequency crowding issues. This may be achieved by minimizing the spectral energy of the control pulses across crosstalk-induced transitions in addition to the $ef$-transition, while using a cutoff frequency to regularize the pulse shape. Even though HD DRAG and FAST DRAG provide an approximately equal leakage error in the studied single-qubit system, we foresee that the flexibility of FAST DRAG may be beneficial to appropriately adjust the relative strength and width of spectral suppression in the case of multiple undesired transitions as discussed in Appendix F.
The proposed methods may also find applications in the optimization of two-qubit gate control pulses based on microwave pulses [67, 6] or flux pulses [5, 4]. As an example, architectures based on the cross-resonance effect [68, 69, 29] or the recently demonstrated fluxonium-transmon-fluxonium scheme [6] rely on microwave-activated two-qubit gates that suffer from frequency crowding. Hence, the analytical pulse shaping methods developed in this work may help to suppress undesired transitions improving the microwave-activated two-qubit gates, and potentially speeding up the calibration compared to, e.g., the reinforcement learning approach employed in Ref. [6]. Furthermore, we expect that combining the proposed methods with experimental closed-loop optimization may result in further improvements to the fidelity of fast quantum logic gates in systems with high coherence. Namely, the proposed methods provide good initial parameter guesses and a natural basis for the optimization, potentially reducing the required number of iterations and wall clock time significantly.
@@
Data availability
Data supporting the findings of this article is available at: https://doi.org/10.5281/zenodo.13365033. The dataset in Zenodo also includes a Jupyter notebook that provides Python helper functions for evaluating and plotting the FAST DRAG and HD DRAG pulse shapes defined by Eqs. (2), (5), (6), and (7).
@@ Acknowledgements.
We thank the whole staff at IQM Quantum Computers for their support. Especially, we acknowledge Lucas Ortega, Roope Kokkoniemi and Matthew Sarsby for supporting the maintenance and improvement of the experimental setup; Tuure Orell, Jani Tuorila, and Hao Hsu for simulation-related discussions; and Attila Geresdi on discussions related to single-qubit gates. E.H. thanks the Finnish Foundation for Technology Promotion (grant No. 9230) and Jenny and Antti Wihuri Foundation (grant No. 00230115) for funding. We acknowledge the provision of facilities and technical support by Aalto University at OtaNano - Micronova Nanofabrication Center and LTL infrastructure. All authors declare that IQM has filed a patent application regarding the new methods for shaping the frequency spectrum of control pulses having the following inventors: Eric Hyyppä, Antti Vepsäläinen and Johannes Heinsoo. E.H. developed the concept and theory for FAST DRAG and HD DRAG. E.H., A.V., and J.H. planned the simulations and the experiments. E.H. conducted the simulations with support from M.P. and S.I.. E.H. conducted the experiments and analyzed the measurement data with support from J.L., F.M., S.O., and C.F.C. on measurement code, F.M. on experiment setup, and B.T. on experiment software. A.L. and C.O.-K designed the sample. W.L. fabricated the device. The manuscript was written by E.H. with support from A.V., and J.H.. All authors commented on the manuscript. A.V., and J.H. supervised the work.
@@ Appendix A Derivation of FAST pulse shapes
where $N$ is the number of basis functions, $\{g_{n}(t)\}$ is the set of basis functions defined to be 0 outside of the interval $[0,t_{\mathrm{p}}]$, and $\{c_{n}\}$ are the corresponding coefficients. In this work, we have chosen the sum of basis functions to represent a Fourier cosine series parametrized such that the control envelope $\Omega_{\mathrm{I}}:\mathbb{R}\rightarrow\mathbb{R}$ and its derivative have no discontinuities, thus reducing slow sinc-type decay of the pulse spectrum. Thus, we write the control envelope $\Omega_{\mathrm{I}}(t)$ as
$\Omega_{\mathrm{I}}(t)=A\times\sum_{n=1}^{N}c_{n}[1-\cos(2\pi nt/t_{\mathrm{p}})]\Pi(t/t_{\mathrm{p}}-1/2),$ (9)
where $\Pi(x)$ is the rectangular window function between $x\in[-1/2,1/2]$, which sets the pulse envelope to 0 outside of $[0,t_{\mathrm{p}}]$.
@@
$=\int_{0}^{t_{\mathrm{p}}}[1-\cos(2\pi nt/t_{\mathrm{p}})]\mathrm{e}^{-\mathrm{i}2\pi tf}\mathrm{d}t$
@@
where $\tilde{\bm{c}}=(\bm{c}^{T},\mu)^{T}$ is the coefficient vector extended by the Lagrangian multiplier, $\tilde{\bm{b}}=(\bm{0}^{T},\theta/t_{\mathrm{p}})^{T}\in\mathbb{R}^{(N+1)\times 1}$ with $\bm{0}=(0,\dots,0)^{T}\in\mathbb{R}^{N\times 1}$, and the matrix $\tilde{\bm{A}}\in\mathbb{R}^{(N+1)\times(N+1)}$ is defined as
@@
Subsequently, we briefly discuss some useful properties of the FAST method. See also Appendix F for guidelines on the selection of the hyperparameters. First, we emphasize that the FAST method enables the construction of pulses with a given duration. Thus, the method differs from filter-based solutions, such as the use of band-pass or low-pass filters [18], which may filter out specific frequency intervals from the spectrum of a control pulse but which also result in pulse distortions, potentially significantly changing the pulse duration. As already discussed in Sec. II.2, the FAST method enables one to control the width and strength of the spectral suppression. Thus, a combination of FAST and DRAG may be helpful to reach a lower leakage error for fast logic gates compared to conventional DRAG pulses since the spectral suppression provided by conventional DRAG pulses may not be sufficient, e.g., in the presence of multiple leakage transitions or frequency shifts occurring during the gate. However, one should bear in mind that the suppression of several frequency intervals using the FAST method may lead to increased spectral power at non-suppressed frequencies.
@@
Finally, we note that other choices for the basis functions $\{g_{n}(t)\}$ beyond the cosine series are also possible and the coefficients $\{c_{n}\}$ of the basis functions can still be solved using Eq. (19). The expression for $\hat{g}_{n}(f)$ needs to be analytically evaluated for the new set of basis functions or numerically computed across the undesired frequency intervals, which may, however, increase the computational cost of solving $\{c_{n}\}$.
@@ Appendix B Higher-derivative DRAG
$\Omega_{\mathrm{Q}}(t)$ $=-\frac{\beta}{\alpha}\dot{\Omega}_{\mathrm{I}},$ (22)
@@
The formulation of HD DRAG in Eqs. (21) and (22) decouples the in-phase parameters $\{\beta_{2n}\}_{n=1}^{K}$ and associated suppressed frequencies $\{f_{j}\}_{j=1}^{K}$ from the DRAG coefficient $\beta$, which simplifies the calibration experiments especially for the case of $K=1$ considered in Sec. II.3 and also enables one to tune $\beta$ to avoid phase errors if desired (see Sec. II.1). By solving both $\{\beta_{2n}\}_{n=1}^{K}$ and $\beta$ to suppress a given undesired baseband frequency $f_{\mathrm{s}}$, a stronger and wider spectral suppression can be obtained than with conventional DRAG pulses. Furthermore, the in-phase parameters $\{\beta_{2n}\}_{n=1}^{K}$ do not affect the effective drive frequency (unlike $\beta$), which can be beneficial for calibration experiments in some scenarios.
@@ Appendix C Single-qubit gate simulations
In this section, we describe the methods used to simulate the gate error and leakage as illustrated in Fig. 3, and in Supplementary Figs. 12--15. For the simulations, we model the bare transmon using an anharmonic Duffing oscillator Hamiltonian as
@@
For a given control pulse shape and gate duration, we set the drive frequency equal to the qubit frequency and calibrate the pulse parameters $\{A,\beta,\varphi_{z}\}$ for a $R_{X}(\pi/2)$-gate based on DRAG-P or DRAG-L using a simulated calibration sequence similar to the one used in the experiments and described in Appendix G. As a minor difference, the leakage per gate needed for calibrating $\beta$ in the case of DRAG-L is estimated based on a single simulated $R_{X}(\pi/2)$-gate instead of using a leakage RB sequence.
@@ Appendix D Experimental setup
To implement single-qubit rotations, the qubits are capacitively coupled to individual drive lines that are connected to room-temperature electronics via 50 dB of total nominal attenuation separated across different temperature stages of the cryostat. At the qubit frequency, the total room-temperature attenuation in the cryostat was characterized to be approximately 67 dB. At room temperature, the single-qubit control pulses are generated at an intermediate frequency using a channel pair of a HDAWG (Zurich Instruments) with a sampling rate of 2.4 GSa/s and a peak-to-peak maximum amplitude range of 3 V or 4 V enabled by an internal amplifier, which limits the analog bandwidth to 300 MHz. To attain a sample-level granularity for the gate duration and facilitate envelope predistortion discussed in Appendix H, we evaluate the $I$ and $Q$ envelope waveforms of each gate sequence on a computer accounting for the Virtual Z-rotations, after which we send these long waveforms to the HDAWG. The intermediate-frequency control pulses are upconverted to the qubit frequency using an IQ mixer and an in-house-built local oscillator with an integrated root-mean-squared phase jitter below 4 mrad from 1 kHz to 10 MHz.
@@
For qubit state readout, the qubits are capacitively coupled to individual readout structures consisting of a readout resonator and a Purcell filter with both of the readout structures further coupled to a common probe transmission line. We generate the readout pulses at an intermediate frequency using an UHFQA (Zurich Instruments) and upconvert the pulses with an IQ mixer and a single SGS100A microwave generator (Rohde & Schwarz) that is used both for the up-conversion and down-conversion of the readout signal. On its way to the sample, the readout signal is passed through 60 dB of nominal attenuation. The readout signal transmitted through the probe line is amplified using an high-electron-mobility transistor (HEMT) at the 4 K stage with further amplification applied at room temperature. The amplified readout signal is downconverted back to an intermediate frequency and subsequently digitized using the UHFQA. Note that our setup does not include a quantum-limited parametric amplifier despite which we reach a decent two-state readout assignment fidelity of up to $\sim 95$%.
@@ Appendix E Three-state readout
In this section, we briefly describe our approach for performing three-state readout on the qubit used for the single-qubit gate studies. At the flux-insensitive operation point, we measure a dispersive shift of $\chi/(2\pi)=4.0$ MHz. For the readout pulse, we use a readout frequency of 6.146 GHz and a relatively long pulse duration of 1.5 $\mu$s due to the absence of a quantum-limited amplifier in the readout chain. For the integration of the acquired readout signal, we use an integration time of 1.5 $\mu$s and set all integration weights to unity. To calibrate the classification boundaries for three-state readout, we repeatedly prepare the qubit in each of the states $|0\rangle$, $|1\rangle$, and $|2\rangle$, and record $326782^{15}=$ 32678 single-shot IQ voltages for each state, as shown in Fig. 7. For each prepared state, the mean of the acquired IQ voltages is computed and for new readout events, the qubit is assigned to the state with the smallest distance between the recorded single-shot IQ voltage and the state mean. The resulting readout assignment matrix is given by
@@
Appendix F Hyperparameter selection for FAST DRAG and HD DRAG
In this section, we provide guidelines for selecting and calibrating the hyperparameters for FAST DRAG ($N$, $\{w_{j}\}$, $\{[f_{\mathrm{l},j},f_{\mathrm{h},j}]\}$) and HD DRAG ($\{\beta_{2n}\}$). We first consider a scenario where the $ef$-transition is the dominant source of leakage as in this work. Subsequently, we briefly cover the expected modifications that are needed to mitigate multiple leakage transitions, e.g., in simultaneous single-qubit gates suffering from crosstalk.
To mitigate leakage to the $f$-level, the hyperparameters of the FAST method can be selected using the following physics-based heuristic algorithm:
  1. 1.
    Parametrize the weight ratio as $w_{\mathrm{ef}}=w_{1}/w_{2}$ and the frequency intervals to be suppressed as $[f_{\mathrm{l},1},f_{\mathrm{h},1}]=[f_{\mathrm{l},\mathrm{ef}},f_{\mathrm{h},\mathrm{ef}}]$ and $[f_{\mathrm{l},2},f_{\mathrm{h},2}]=[f_{\mathrm{c}},\infty]$. Now, the hyperparameters to be determined are $\{f_{\mathrm{l},\mathrm{ef}},f_{\mathrm{h},\mathrm{ef}},f_{\mathrm{c}},w_{\mathrm{ef}},N\}$.
  2. 2.
    Measure the anharmonicity $\alpha/(2\pi)$ of the qubit in Hz.
  3. 3.
    Set the center of the first frequency interval to the measured anharmonicity and the width, e.g., as $0.1\times|\alpha|/(2\pi)$ ($\sim 20$ MHz for typical transmon anharmonicities) resulting in $f_{\mathrm{l},\mathrm{ef}}=0.95\times|\alpha|/(2\pi)$ and $f_{\mathrm{h},\mathrm{ef}}=1.05\times|\alpha|/(2\pi)$.
  4. 4.
    Set the cutoff frequency as $f_{\mathrm{c}}=2|\alpha|/(2\pi)$ based on the measured anharmonicity.
  5. 5.
    Set the weight as $w_{\mathrm{ef}}=5$ for DRAG-L and $w_{\mathrm{ef}}=100$ for DRAG-P similarly to the experiments in this work.
  6. 6.
    The number of Fourier terms can be set, e.g., as $N=4$.
  7. 7.
    To implement single-qubit gates, use Eqs. (2) and (5) to obtain the pulse shape for the selected hyperparameters
For HD DRAG, the procedure is simpler as the only additional parameter is $\beta_{2}$, which can be set as $\beta_{2}=1/\alpha^{2}$ to mitigate $f$-level leakage. According to simulations shown in Fig. 8(a), the above guidelines are expected to outperform Cosine DRAG-L at least for typical transmon anharmonicities $|\alpha|/(2\pi)\in(150,300)$ MHz across fast gate durations $t_{\mathrm{g}}|\alpha|/(2\pi)\in(1,2)$, for which Cosine DRAG-L is not sufficient to suppress leakage errors. Note that the hyperparameters obtained with these guidelines are very close to the parameters used for the experiments in Sec. III.
Figure 8: Hyperparameter selection for FAST DRAG. (a) Simulated leakage per gate $L_{\mathrm{g}}$ as a function of the normalized gate duration $t_{\mathrm{g}}|\alpha|/(2\pi)$ for different anharmonicities $\alpha/(2\pi)\in\{-100,-160,-212,-300\}$ MHz (shades of blue) when selecting the hyperparameters of FAST DRAG-L using the heuristic algorithm of Appendix F. For comparison, we show the leakage of Cosine DRAG-L (red line) for $\alpha/(2\pi)=-212$ MHz. In the simulations, we assume $T_{1}=35$ $\mu$s, $T_{\varphi}=40$ $\mu$s, and $\bar{n}=0.02$ as explained in Appendix C. (b) Optimal value for the center of the suppressed frequency interval $(f_{\mathrm{l},\mathrm{ef}}+f_{\mathrm{h},\mathrm{ef}})/2$ (dark blue) to minimize simulated leakage per gate $L_{\mathrm{g}}$ as a function of the gate duration for FAST DRAG-L when selecting other parameters based on the heuristic algorithm of Appendix F. The colored area shows values of $(f_{\mathrm{l},\mathrm{ef}}+f_{\mathrm{h},\mathrm{ef}})/2$ yielding a leakage per gate at most twice the optimal value, while the dashed line corresponds to $|\alpha|/(2\pi)$. For each gate duration, the DRAG coefficient $\beta$ is calibrated assuming $(f_{\mathrm{l},\mathrm{ef}}+f_{\mathrm{h},\mathrm{ef}})/2=|\alpha|/(2\pi)$ and kept fixed while sweeping $(f_{\mathrm{l},\mathrm{ef}}+f_{\mathrm{h},\mathrm{ef}})/2$.
The above heuristic guidelines are based on the following reasoning. In the experiments and the simulations, the optimal center frequency $(f_{\mathrm{l},\mathrm{ef}}+f_{\mathrm{h},\mathrm{ef}})/2$ is observed to approximately equal the anharmonicity $|\alpha|/(2\pi)$, with the minimum of leakage being relatively broad, as shown in Figs. 13(e) and 14(e) in Appendix J. By selecting $(f_{\mathrm{l},\mathrm{ef}}+f_{\mathrm{h},\mathrm{ef}})/2=|\alpha|/(2\pi)$, one is expected to obtain a leakage rate that is typically within a factor of two of the optimal leakage rate, see Fig. 8(b). The choice of the cutoff frequency $f_{\mathrm{c}}=2|\alpha|/(2\pi)$ provides a reasonable balance between leakage suppression and regularization of the pulse shape, which ensures a low leakage error for fast gates while limiting the peak amplitude compared to Cosine DRAG-L pulses (see Fig. 3(f)). When it comes to the weight $w_{\mathrm{ef}}$ and the interval width $f_{\mathrm{h},\mathrm{ef}}-f_{\mathrm{l},\mathrm{ef}}$, we observe that the resulting leakage error for FAST DRAG-L pulses is largely insensitive to the exact value of these parameters both in simulations and experiments as long as the parameters have the correct order of magnitude, see Figs. 14(d) and (f) in Appendix J. However, we observe that too aggressive spectral suppression with an excessively high value of $w_{\mathrm{ef}}$ may lead to increased leakage errors with the DRAG-L calibration, as shown in Fig. 14(d).
Figure 9: Regularizing effect of the cutoff frequency on the Fourier coefficients in FAST DRAG pulses. (a) Absolute value of the Fourier coefficients $|c_{n}|$ as a function of pulse duration for a FAST DRAG pulse with $N=6$ and a single suppressed frequency interval $[f_{\mathrm{ef},\mathrm{l}},f_{\mathrm{ef},\mathrm{h}}]=[194,214]$ MHz, i.e., $f_{\mathrm{c}}\to\infty$ or $w_{\mathrm{ef}}\to\infty$. (b) Critical number of basis functions $N_{\mathrm{crit}}$ as a function of pulse duration for a FAST DRAG pulse with $f_{\mathrm{c}}\to\infty$ or $w_{\mathrm{ef}}\to\infty$. (c) Same as (a) but using a second frequency interval to specify a cutoff frequency $f_{\mathrm{c}}=450$ MHz and a weight $w_{\mathrm{ef}}=100$ controlling the relative suppression across the two frequency intervals.
Furthermore, there is some freedom in the selection of the number of Fourier terms since already $N=2$ provides a low leakage rate for FAST DRAG-L in the studied system with one dominant leakage transition, as shown in Fig. 14(b). However, a higher number of basis functions $N\geq 3$ may be beneficial to suppress the in-phase spectrum across more than one leakage transitions or to control the strength of the spectral suppression. This is exemplified in Fig. 13(b) demonstrating that the leakage errors caused by FAST DRAG-P pulses are reduced by increasing the number of basis functions, which allows a stronger spectral suppression around the $ef$-transition for fast gates. Importantly, the use of a finite cutoff frequency $f_{\mathrm{c}}$ and weight $w_{\mathrm{ef}}$ render the pulse shape and the leakage error less sensitive to the choice of $N$. For $f_{\mathrm{c}}\to\infty$ or $w_{\mathrm{ef}}\to\infty$, there exists a critical value $N_{\mathrm{crit}}$ below which the choice of $N$ naturally limits the bandwidth and peak power but above which the Fourier coefficients $\{c_{n}\}$ attain high absolute values leading to a high peak power and bandwidth of the control pulse. Here, we define the critical value $N_{\mathrm{crit}}$ as the largest value of $N$ ensuring that the highest-frequency component remains bounded according to the condition $|c_{N}|<\theta/t_{\mathrm{p}}$ in the case of $f_{\mathrm{c}}\to\infty$ or $w_{\mathrm{ef}}\to\infty$, where $\theta$ is the rotation angle of Eq. (4) As the pulse duration is reduced, the Fourier coefficients of high-frequency components increase first as illustrated in Fig. 9(a), thus implying that the critical value $N_{\mathrm{crit}}$ is reduced for short gate durations, as shown in Fig. 9(b). For fast gates with $t_{\mathrm{p}}|\alpha|/(2\pi)\in(1,2)$, the number of basis functions would need to be limited to $N=2$ to ensure a reasonable bandwidth and peak power in the case of $f_{\mathrm{c}}\to\infty$ or $w_{\mathrm{ef}}\to\infty$. For these reasons, we favor the use of a finite cutoff frequency and weight, which lead to bounded Fourier coefficients across gate durations of interest, see Fig. 9(c). Thus, the cutoff frequency allows us to regularize the pulse shape, while still benefiting from the stronger and more flexible spectral suppression and potentially reduced leakage errors obtained with increasing $N$.
We note that further performance improvements compared to the heuristic algorithm may be achieved by performing closed-loop optimization of the hyperparameters as a part of the single-qubit gate calibration. The parameter values predicted by the heuristic algorithm can be used as educated initial values to enhance convergence.
In the presence of multiple leakage transitions, e.g., in simultaneous single-qubit gates, the heuristic algorithm for FAST DRAG needs to be extended. Similarly to the single-qubit case, one should measure the frequencies of potential undesired transitions, and estimate the coupling strength to each of these transitions. Subsequently, one can choose to suppress frequency intervals corresponding to the $ef$-transition and the dominant crosstalk-induced transitions, while reserving an additional frequency interval to regularize the pulse shape. A good starting point for FAST DRAG is to set the center frequencies of the intervals equal to the frequencies of the undesired transitions. To enable strong spectral suppression across more than one frequency intervals, one should also use $N\geq 3$ basis functions. Furthermore, one should be aware that aggressive spectral suppression across multiple frequency intervals leads to increased spectral power across non-suppressed frequencies and a wider bandwidth, which may have unintended consequences in systems with high crosstalk or with many transitions to avoid. Thus, further research is needed to design reliable heuristics for setting the weights and widths of the frequency intervals using the measured frequencies and the associated coupling strengths. With HD DRAG, the suppressed baseband frequencies $\{f_{j}\}$ can be directly set to the undesired transition frequencies, after which the parameters $\{\beta_{2n}\}$ can be solved from Eq. (25). As emphasized in Sec. II.3, HD DRAG cannot control the relative strength of spectral suppression between the different frequency intervals in contrast to FAST DRAG that provides more flexibility, which is expected to enable an enhanced performance in systems with multiple leakage transitions. As the multi-qubit system is more challenging and the required number of parameters is higher compared to the single-qubit system, closed-loop optimization may be needed to reach close-to-optimal performance. However, physics-based heuristics are expected to help in finding good initial values for the optimization.
Appendix G Single-qubit gate calibration